# Visual Perception Engine: Fast and Flexible Multi-Head Inference for Robotic Vision Tasks

### v1

@@ Abstract
Deploying multiple machine learning models on resource-constrained robotic platforms for different perception tasks often results in redundant computations, large memory footprints, and complex integration challenges. In response, this work presents Visual Perception Engine (VPEngine), a modular framework designed to enable efficient GPU usage for visual multitasking while maintaining extensibility and developer accessibility. Our framework architecture leverages a shared foundation model backbone that extracts image representations, which are efficiently shared, without any unnecessary GPU-CPU memory transfers, across multiple specialized task-specific model heads running in parallel. This design eliminates the computational redundancy inherent in feature extraction component when deploying traditional sequential models while enabling dynamic task prioritization based on application demands. We demonstrate our framework’s capabilities through an example implementation using DINOv2 as the foundation model with multiple task (depth, object detection and semantic segmentation) heads, achieving up to 3x speedup compared to sequential execution. Building on CUDA Multi-Process Service (MPS), VPEngine offers efficient GPU utilization and maintains a constant memory footprint while allowing per-task inference frequencies to be adjusted dynamically during runtime. The framework is written in Python and is open source with ROS2 C++ (Humble) bindings for ease of use by the robotics community across diverse robotic platforms. Our example implementation demonstrates end-to-end real-time performance at $\geq$50 Hz on NVIDIA Jetson Orin AGX for TensorRT optimized models.
@@ I Introduction
Figure 1: High level architectural overview of the VPEngine framework.
To address these limitations, we introduce Visual Perception Engine (VPEngine), a framework specifically designed to harness the power of visual foundation models for robotic multitasking efficiently. VPEngine uses highly optimized TensorRT models [^1] However, it supports native PyTorch as well, since it is not always feasible to convert a model to TensorRT, running in separate processes, which combined with CUDA MPS results in a very fast inference and high GPU utilization. Furthermore, we devised custom Inter-Process Communication structures using CUDA API to enable by reference GPU memory transfers across processes, which are not supported on Jetson devices. As a consequence, the latency of sharing image representations between the backbone and model heads is negligible. The framework is written in Python, heavily leverages abstract classes and is highly modularized which enables fast prototyping and fosters extendability. A high-level overview of the system is shown in Figure 1.
@@ II Design Requirements
To address the challenges of efficient visual multitasking in resource-constrained robotic systems, we establish the following design requirements that guide VPEngines’s architecture and implementation:
  • R1: Fast Inference - Latency of the perception loop can be the upper bound on the robots reaction speed to a change in the environment. Furthermore, if desired model frequencies are fixed, inference efficiency allows for more GPU time available for other tasks.
@@ III System architecture
Figure 2: Alternative High level architectural overview of the VPEngine framework.
@@ III-A4 Module separation
Each model component executes in a separate process to maximize GPU utilization through CUDA MPS, addressing both R1 and R3. To utilize CUDA MPS, a standard MPS server is configured on the deployment device, after which all CUDA-enabled processes automatically leverage shared GPU resources. Additional benefit of using CUDA MPS is that it allocates only one CUDA context for all processes[8].
@@ IV Example implementation
In our example implementation, we use three model heads: DepthAnythingV2 (DAV2) [9] responsible for monocular depth estimation (we used optimized code from [10]) compiled with TensorRT, linear layer for semantic segmentation [4] compiled with TensorRT, and a custom FasterRCNN head for 2-class object detection. For foundation model we use ViT-S version of DINOv2 [4] to get rich visual representations. We configured it to output final features along with representations computed at three evenly spaced intermediate layers into the middle buffer, since they are required by DAV2 model head.
We test our example implementation on a stream of 30000 images of $1920\times 1080$ resolution and it maintains [^2] Due to freshness-oriented design and stochastic nature of the engine minimal losses are unavoidable, especially when using models not compiled with TensorRT. However, even then they account for $\approx 0.02\%$ of all frames. a 30 Hz throughput with median latencies of 2020ms for depth estimation, 1818ms for semantic segmentation, and 6969ms for object detection on a Nvidia Orin AGX. It is important to note that the object detection head is not compiled with TensorRT, hence larger latency. Throughout the test, GPU memory usage was constant at around 1.51.5GB. Using VPEngine, the combined number of parameters of our example implementation is 27M, which 62% less than if we used three independent models, one for each task (71M parameters total).
V Benchmark
@@
Regarding memory, the full DepthAnythingV2 model uses an additional 96MB of GPU memory for each new task, the Sequential Heads baseline uses 48MB, and VPEngine uses 214MB per task.
@@
This experiment quantifies the advantage of VPEngine’s ability to run multiple model heads in parallel. We compared our framework against a “Sequential Object Detection Heads” baseline, which uses a single backbone but runs multiple PyTorch-based object detection heads one after another. Unlike the TensorRT kernels in the previous experiment, these PyTorch heads do not fully utilize the GPU.
@@ VI Limitations
The most significant limitation of our framework is the use of multiple processes, leading to higher CPU usage. This also results in slightly higher GPU memory usage [^4] In our benchmarks we are using relatively small models, hence extra 295295MB seems large in comparison, but please note that many models such as DINOv2-Giant use order of magnitude more memory (4.4GB in this case)..
Secondly, different modules are running independently, which leads to overall stochastic nature of VPEngine. Implementing coordination between different processes, could potentially result in even higher GPU usage. We leave this to future work.
@@ Appendix C CUDA MPS Ablation Results
We ablate the gain from using CUDA MPS by using the VPEngine setup from Section LABEL:sec:parallel_design and measuring the speed with CUDA MPS first disabled and then enables. Results are in Table.I

### v2

@@ Abstract
Deploying multiple machine learning models on resource-constrained robotic platforms for different perception tasks often results in redundant computations and complex integration challenges. In response, this work presents Visual Perception Engine (VPEngine), a modular framework designed to enable efficient GPU usage for visual multitasking while maintaining extensibility and developer accessibility. Our framework architecture leverages a shared foundation model backbone that extracts image representations, which are efficiently shared, without any unnecessary GPU-CPU memory transfers, across multiple specialized task-specific model heads running in parallel. This design eliminates the computational redundancy inherent in feature extraction component when deploying traditional sequential models while enabling dynamic task prioritization based on application demands. We demonstrate our framework’s capabilities through an example implementation using DINOv2 as the foundation model with multiple task (depth, object detection and semantic segmentation) heads, achieving up to 3x speedup compared to sequential execution. Building on CUDA Multi-Process Service (MPS), VPEngine offers efficient GPU utilization and maintains a constant memory footprint while allowing per-task inference frequencies to be adjusted dynamically during runtime. The framework is written in Python and is open source with ROS2 C++ (Humble) bindings for ease of use by the robotics community across diverse robotic platforms. Our example implementation demonstrates end-to-end real-time performance at $\geq$50 Hz on NVIDIA Jetson Orin AGX for TensorRT optimized models.
@@ I Introduction
Figure 1: Example implementation use case of the proposed Visual Perception Engine (VPE) for three downstream robotic vision tasks. VPE enables without copying sharing of visual foundation features from TensorRT-optimized backbone with multiple independently TensorRT-optimized (depth estimation and semantic segmentation) or PyTorch (object detection) task heads. The CPU and GPU tasks are represented by white and black boxes, respectively.
To address these limitations, we introduce Visual Perception Engine (VPEngine), a framework specifically designed to harness the power of visual foundation models for robotic multitasking efficiently. VPEngine uses highly optimized TensorRT models [^1] However, it supports native PyTorch as well, since it is not always feasible to convert a model to TensorRT., running in separate processes, which combined with CUDA MPS results in a very fast inference and high GPU utilization. Furthermore, we devised custom Inter-Process Communication structures using CUDA API to enable by reference GPU memory transfers across processes, which are not supported on Jetson devices. As a consequence, the latency of sharing image representations between the backbone and model heads is negligible. The open-source  [^2] https://github.com/nasa-jpl/visual-perception-engine framework is written in Python, heavily leverages abstract classes and is highly modularized which enables fast prototyping and fosters extendability. A high-level overview of the system is shown in Figure 1.
@@ II Design Requirements
To address the challenges of efficient visual multitasking in resource-constrained robotic systems, we establish the following design requirements that guide VPEngines’s architecture:
R1: Fast Inference - Perception loop latency can be the upper bound on the robot’s reaction speed to environment changes. Furthermore, if desired model inference frequencies are fixed, efficiency allows for GPU availability for other tasks.
@@ III System architecture
Figure 2: Architecture overview of the VPE showing interaction between CPU and GPU computations. The figure also highlights preservation of visual foundation model feature in GPU memory and efficiently sharing with with Pytorch modules and TensorRT optimized kernels minimizing the number of copies.
@@ III-A4 Module separation
Each model component executes in a separate process to maximize GPU utilization through CUDA MPS, addressing both R1 and R3. To utilize CUDA MPS, a standard CUDA MPS server is configured on the deployment device, after which all CUDA-enabled processes automatically leverage shared GPU resources. Additional benefit of using CUDA MPS is that it allocates only one CUDA context for all processes[8].
@@ IV Example implementation
In our example implementation, we use three model heads: DepthAnythingV2 (DAV2) [9] responsible for monocular depth estimation (we used optimized code from [10]) compiled with TensorRT, linear layer for semantic segmentation [4] compiled with TensorRT, and a custom FasterRCNN head object detection. To get vision foundation model representations we use ViT-S version of DINOv2 [4]. We configure it to output final features along with representations computed at three evenly spaced intermediate layers into the middle buffer, since they are required by DAV2 model head.
We test our example implementation on a stream of 30000 images of $1920\times 1080$ resolution, and it maintains  [^3] Due to freshness-oriented design and stochastic nature of the engine minimal losses are unavoidable, especially when using models not compiled with TensorRT. However, even then they account for $\approx 0.02\%$ of all frames. a 30 Hz throughput with median inference times of 2020ms, 1818ms and 3232ms for the depth estimation, semantic segmentation and object detection heads, respectively, on a Nvidia Orin AGX 64GB. It is important to note that the object detection head is not compiled with TensorRT, hence the larger inference time. Throughout the test, GPU memory usage was constant at around 1.51.5GB. Using VPEngine, the combined number of parameters of our example implementation is 27M [^4] Using our custom object detection head and the others provided with [4]., which 69% lower than if we used three independent models  [^5] Using DINOv2 models for semantic segmentation and depth estimation, and FasterRCNN implementation in Pytorch.
V Benchmarks
@@ VI Limitations
A limitation of our framework is the use of multiple processes, leading to higher CPU usage. This also results in slightly higher GPU memory usage [^7] In our benchmarks we are using relatively small models, hence extra 295295MB seems large in comparison, but please note that many models such as DINOv2-Giant use order of magnitude more memory (4.4GB in this case).. Additionally, different modules are running independently, which leads to overall stochastic nature of VPEngine. Implementing coordination between different processes, could potentially result in even higher GPU usage. We leave this to future work.
@@ Appendix C CUDA MPS Ablation Results
We ablate the gain from using CUDA MPS by using the VPEngine setup from Section V-B and measuring the speed with CUDA MPS first disabled and then enables. Results are in Table.I